500gm of salt solution has 30% salt in it.how much salt should be added to make the salt 50% in the solution?
Answer by stanbon(75887)

500gm of salt solution has 30% salt in it.how much salt should be added to make the salt 50% in the solution?
-------
Equation:
salt + salt = salt
0.30*500 + 1*x = 0.50(500+x)
---------
30*500 + 100x = 50*500 + 50x
---------
50x = 20*500
x = 200 grams (amt. of pure salt to add)
-----------------
Cheers,
Stan H.
---------------